OpenStudy (anonymous):

Medal!!! The temperature decreased overnight by 15.2°C to a temperature of –12.8°C. Suppose t represents the temperature before it decreased. Which equation can be used to represent this situation?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

A. t – 15.2 = –12.8 B. t + 12.8 = –15.2 C. t – 12.8 = 15.2 D. t – (–15.2) = –12.8

OpenStudy (anonymous):

Do you know what it is

OpenStudy (igreen):

It decreased by 15.2, and the result was -12.8. Therefore, t - 15.2 = -12.8

OpenStudy (anonymous):

Thanks!

OpenStudy (igreen):

No problem!

OpenStudy (anonymous):

1 more?

OpenStudy (igreen):

Sure.